A 10mV, 2KHz sinusoidal signal is applied to the input of an op-amp integrator for which R=100 kΩ and C=1 μF. Find the output voltage.


1
Expert's answer
2021-05-31T06:18:38-0400

"V_o=-\frac{1}{RC}\int v_idt"

"v_{in}= v_msin \omega t=10 sin 2\pi ft=10 sin4000 \pi t \space mV"

So, "\frac{-1}{100*10^3*1*10^{-6}}\int 10 sin 4000 \pi tdt \space mV = \frac{0.025}{\pi}cos 4000 \pi t \space \space mV"